Splish Company produces one product, a putter called GO-Putter. Splish uses a standard cost system and determines that it should take one hour of direct labor to produce one GO-Putter. The normal production capacity for this putter is 145,000 units per year. The total budgeted overhead at normal capacity is $942,500 comprised of $290,000 of variable costs and $652,500 of fixed costs. Splish applies overhead on the basis of direct labor hours.

During the current year, Splish produced 81,600 putters, worked 87,300 direct labor hours, and incurred variable overhead costs of $159,120 and fixed overhead costs of $471,400.

Compute the predetermined variable overhead rate and the predetermined fixed overhead rate. (Round answers to 2 decimal places, e.g. 2.75.)

Variable

Fixed

Predetermined Overhead Rate $

$

LINK TO TEXT

Compute the applied overhead for Splish for the year.
Overhead Applied $

LINK TO TEXT

Compute the total overhead variance.
Total Overhead Variance $

Neither favorable nor unfavorableUnfavorableFavorable

Click if you would like to Show Work for this question:

Open Show Work
Answer #1
1
Predetermined Overhead Rate :
Variable 2.00 =290000/145000
Fixed 4.50 =652500/145000
2
Standard hours for actual output 81600
X Predetermined Overhead Rate 6.50 =2.00+4.50
Overhead Applied 530400
3
Total Actual overhead 630520 =159120+471400
Less: Overhead Applied 530400
Total Overhead Variance 100120 Unfavorable